OpenStudy (mathmath333):

The letters of the word ALLAHABAD are rearranged to form new words and put in a dictionary. If the dictionary has only these words and one word on every page in alphabetical order then what is the page number on which the letter LABADALAH comes?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ The letters of the word ALLAHABAD are rearranged to form }\hspace{.33em}\\~\\ & \normalsize \text{ new words and put in a dictionary. If the dictionary has only}\hspace{.33em}\\~\\ & \normalsize \text{ these words and one word on every page in alphabetical}\hspace{.33em}\\~\\ & \normalsize \text{ order then what is the page number on which the letter }\hspace{.33em}\\~\\ & \normalsize \text{ LABADALAH comes? }\hspace{.33em}\\~\\ \end{align}}\)

Nnesha (nnesha):

sigh.

Nnesha (nnesha):

is that the original question or did you change some words ? just asking....

OpenStudy (mathmath333):

It is 100% genuine.

Nnesha (nnesha):

okay nvm.it's bank name i guess...

OpenStudy (kropot72):

The total number of pages in the book, with only one new word on each page, is given by: \[\large \frac{9!}{4!2!}=7560\ pages\]

OpenStudy (empty):

I can never tell when there's a new one of these questions, for the longest time I just assumed they were always the same "counting question" from looking to the left haha

OpenStudy (kropot72):

The number of pages where the first letter of each new word is L is given by: 8!/4! = 1680 pages. Therefore the number of the first page where the first letter of the new word is L is 7560 - 1680 = page 5880. Now we need to find the number of pages after the word LAAAABDHL that occur until we reach the word LABADALAH.

OpenStudy (kropot72):

The number of new words beginning with LAA is given by 6!/2 = 360. Therefore the number of the first page with new words beginning with LAB is found from: 5880 + 360 + 1 = page 6241. The number of new words where the first letters are LABAA is given by 4! = 24. Therefore the number of the first page with new words beginning with LABAD is found from: 6241 + 24 +1 = page 6266.

OpenStudy (kropot72):

Page 6267 has the word LABADAAHL. Page 6268 has the word LABADAALH. PAGE 6269 has the word LABADAHAL.

OpenStudy (beginnersmind):

@Nnesha "okay nvm.it's bank name i guess..." It's a city in India with a population of over a million.

OpenStudy (mathmath333):

thnks

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @beginnersmind @Nnesha "okay nvm.it's bank name i guess..." It's a city in India with a population of over a million. \(\color{blue}{\text{End of Quote}}\) *i guess* i just read the heading https://gyazo.com/8676ed0390348c4c7db93901200ffe25 so i thought it's a bank XD according to wikipedia it supposed to be `Ilahabad` nd then British changed to `Allahabad` dun't know why..... it pronounced like `Aallahbaad` not *bad*

Nnesha (nnesha):

sorry mathmath333 ,-, ...

OpenStudy (mathmath333):

I 'the monk" forgives u nesha

OpenStudy (kropot72):

You're welcome :)

Nnesha (nnesha):

thank u :P it's *nnesha :D :D